Find the surface area of a soup can that has a diameter of 3 inches and a height of 5 inches.
Found 2 solutions by rfer, MathTherapy:
Answer by rfer(16322)

the circumference is 3.14 in
the heigth is 5in
two ends that have 3 in diameter
2(3.14(1.50^2))=14.13 sq in
a=lh
a=(3.14)(5)=15.7 sq in
total area=15.7+14.13=29.83 sq in

Answer by MathTherapy(10552)

The formula for the surface area of a cylinder, which a can is, is:
(pi)*d*h +2(pi)*r^2

You can do the rest, I'm pretty sure.